OpenStudy (anonymous):

find the x-and y- intercepts of the line that passes through the given points

OpenStudy (anonymous):

(-6,4),(3,-5)

OpenStudy (esshotwired):

the easiest way to do this is graph it. but you can also find the equation for it by finding the slope, then plugging it into point slope then solving that to get the slope intercept form of the equation. then make x=0 for the y-intercept and y=0 for x-intercept. but i will give u a link to the graph to make life easier. give me a sec

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

um can you give me an example thats confusing so i turn it into an equation but how

OpenStudy (esshotwired):

first find the slope using this equation: \[\frac{ y ^{2}-y ^{1} }{ x ^{2}-x ^{1} }\] then after u find the slope plug in one of the points(doesnt matter which one) and the slope into the point slope form which is: \[y-y ^{1}=slope(x-x ^{1})\] after the solve to get the equation that looks like this (slope intercept form): \[y=mx+b\]and m=slope

OpenStudy (anonymous):

ok let me solve it and i will answer you

OpenStudy (esshotwired):

alrighty :)

OpenStudy (anonymous):

i got y=-x+4 so x is 6 and y is 4???

OpenStudy (esshotwired):

actually the equation would end up being y=-x-2. tell me ur step so i can find ur problem

OpenStudy (anonymous):

wait how?

OpenStudy (anonymous):

i think i messed up

OpenStudy (esshotwired):

just type in your steps so i can find ur problem and we can fix it

OpenStudy (anonymous):

after i did\[\frac{ y2-y1 }{ x2-x1 }=\frac{ -5-4 }{ 3--6 }=\frac{ -9 }{ 9 }=-1\]

OpenStudy (esshotwired):

u got the slope correct

OpenStudy (anonymous):

then i did y-y1=m(x-x1) y-4=-1(x+6) y=-x+4 ]then i have to find the x- and y- intercept

OpenStudy (anonymous):

i dont know how to do it please help me

OpenStudy (esshotwired):

when you distribute it should look like this:  that would make 6 go back to a negative and also x will be negative to. so then it would be y-4=-x-6 then you would add 4 to both sides and -6+4 is -2